OpenStudy (anonymous):

I Need Guidance :| -2x + y = 1 -4x + y = -1

OpenStudy (texaschic101):

tsolving by substitution or elimination ?

OpenStudy (texaschic101):

typo...solving

OpenStudy (texaschic101):

lets try substitution.. take either equation and try to isolate a variable. It does not matter which one. However, it looks like the easiest one to isolate is the y value. -2x + y = 1 --- add 2x to both sides because what you do on one side of the equal sign, you have to do on the other side. and we would get : y = 2x + 1 now sub 2x + 1 in for y in the other equation -4x + y = -1 -4x + 2x + 1 = -1 -- subtract 1 from both sides -4x + 2x = -1 - 1 -- combine like terms -2x = -2 -- divide both sides by -2 x = 1 now sub 1 in for x in either equation to find y -2x + y = 1 -2(1) + y = 1 -2 + y = 1 y = 2 + 1 y = 3 you can then check your answers by subbing them into either equation to see if they come out equal. -4x + y = -1 -4(1) + 3 = -1 -4 + 3 = -1 -1 = -1 (correct) so x = 1 and y = 3 or (1,3)

OpenStudy (anonymous):

Sorry i didin't get back to you. You explained it very well and i thank you for that.